### v2

@@ 3.1.2 Smart Contract
Calling an SC function intended for blockchain writing involves the execution of a transaction that changes the state of the storage data on the blockchain. In our approach, this transaction is signed with the account’s private key via the Elliptic Curve Digital Signature Algorithm (ECDSA), effectively identifying the signer.
@@ 3.2 Explainability Component
Effective communication is essential for achieving explainability in autonomous robots. Our approach is focused on improving this feature through the use of Natural Language Processing (NLP) technology, enabling non-expert users to interact with agents in a meaningful way.
@@ 4.1 Black Box Recorder
ROS is the most commonly used framework in robotics, consisting of a set of software libraries and tools for building robot applications. However, challenges in meeting the requirements of real-time embedded systems, and the need to improve its multi-platform capabilities led to a major upgrade to ROS 2, whose main change lies in using the Data Distribution Service (DDS) [65].
Nodes constitute the fundamental building block in both ROS and ROS 2 applications. They consist of individual processes, each performing specific tasks and communicating with each other via messages. Data can be moved between nodes using different methods, including topics, services, and actions. Topics are used for message passing, while services provide synchronous request-reply interactions. Actions, on the other hand, are needed for long-running tasks that require feedback.
This work focuses on the navigation functionality. To this end, we have used Nav2, a collection of tools for ROS 2, which is considered the successor of the ROS Navigation Stack. Nav2 provides the robot with the ability to perform complex navigation tasks.
Its main functionalities include planning, control, localization, visualization and other additional features. Furthermore, it uses behavior trees to create customized navigation behavior via orchestrating many independent modular servers, which implies high flexibility in the navigation task and the specification of complex robot behaviors [66]. Different independent task servers, intended to compute a path, control effort, or recovery, are communicated with the behavior tree over a ROS 2 interface, such as an action server or service. The use of behavior trees to generate explanations for robot behavior has been previously successfully evaluated, allowing seeking causal information to answer questions [67].
ROS 2 nodes publish data that other nodes consume through ROS 2 topics. Such data include perception information from sensors or commands to actuators from decision-making nodes. Navigation reproducibility of autonomous robots in ROS 2 involves storing essential data, such as their poses, maps, odometry, velocity, and laser scan readings. To this aim, we have used Rosbag files, a binary file format designed to capture and store data generated during the execution of a ROS system. These files can reproduce the exact conditions and inputs from a previous run, making them particularly valuable for understanding system behavior. The set of topics included in the Rosbag files generated are detailed in Table 1.
@@
To ensure the immutability of generated Rosbag files, Ethereum has been selected as the blockchain framework. Ethereum, an open-source public blockchain, introduces the cryptocurrency Ether and self-executing programs (SCs) for code execution on the blockchain. Ethereum’s open-source nature ensures decentralization and accessibility, allowing participation without needing approval, a feature crucial for maintaining the integrity of recorded data. Additionally, Ethereum’s use of Proof of Stake (PoS) as a consensus protocol enhances energy efficiency and throughput compared to traditional protocols such as Proof of Work (PoW). In addition, Ethereum’s scalability, security and economic efficiency make it a very secure option for ensuring data immutability. The high-level algorithms 1 and 2 depict the main steps in recording and proof verification services.
@@ 4.2 Log Curation Component
The status of the navigation process is checked through the study of the ROS 2 action server NavigateToPose, used for commanding a robot to move to a specified pose. By default, Nav2 guarantees to find the shortest path under all conditions. The /plan topic includes the set of poses computed to reach a specific goal pose. If this initial path changes due to an external circumstance, the Euclidean distance between all the poses included in the pre-computed path will increase. This fact, along with the values obtained through the /scan topic, allows us to determine if the appearance of an obstacle has caused the route modification. These reasonings have been successfully checked in previous authors’ work [68] with the aim of building explanations for autonomous agents based on the use of ROS 2 topics.
@@
6:     // Check a change in the planned path.
7:     if $msgTopic$ == ’/plan’  then
8:         $distance\leftarrow$ getEuclideanDistance($message.poses$)
9:         if $distance>previousDistance$ DISTANCE_THRESHOLD then
10:              $logMsg\leftarrow$ “Planned path has changed when trying to achieve goal pose number $nGoal$"
11:              $previousDistance\leftarrow distance$
12:              if min($last\_scan\_msg.ranges$) $<$ OBSTACLE_DISTANCE_THRESHOLD then
13:                  $logMsg\mathrel{+}=$“ because there was an obstacle."
14:              end if
15:         end if
@@
17:     // Check the navigation status.
18:     if $msgTopic$ == ’/navigate_to_pose/_action/status’  then
19:         $logMsg\leftarrow$ “No navigation is running."
20:         $statusList\leftarrow$ /navigate_to_pose/_action/status
21:         for $status$ in $statusList$ do
22:              $currentGoal\leftarrow statusList.id$
23:              if $currentGoal$ != $previousGoal$ and $currentGoal$ not in $finishedGoals$ then
24:                  $logMsg\leftarrow$ “Navigation to a new goal has started."
25:              else
26:                  $logMsg\leftarrow$ “Navigation to the goal "
27:                  if $status$ = EXECUTING then
28:                       $logMsg\mathrel{+}=$ “is in progress.”
29:                  else if $status$ = SUCCEEDED then
30:                       $logMsg\mathrel{+}=$ “has succeeded."
31:                  else if $status$ = CANCELED then
32:                       $logMsg\mathrel{+}=$ “was cancelled."
33:                  else if $status$ = ABORTED then
34:                       $logMsg\mathrel{+}=$ “has been aborted."
35:                  end if
36:                  if $status$ = SUCCEEDED or $status$ = CANCELED or $status$ = ABORTED  then
37:                       $finishedGoals\cup=currentGoal$
38:                  end if
39:              end if
40:              // Check the executed Behavior Tree nodes and their status.
41:              if $msgTopic$ == ’/behavior_tree_log’  then
42:                  $logMsg\leftarrow$“Node " $message.NodeName$ “performing "$btNodeDesc$“is in status " $message.nodeStatus$
43:              end if
44:              // Gets robot’s position and orientation.
45:              if $msgTopic$ == ’/amcl_pose’  then
46:                  $logMsg\leftarrow$ “Position: " $message.x,message.y$ “Orientation: " $message.z,message.w$
47:              end if
48:              if $msgTopic$ == ’/cmd_vel’  then
49:                  $logMsg\leftarrow$ “Linear velocity: " $message.x,message.y$ “Angular velocity: " $message.z$
50:              end if
51:         end for
52:     end if
53:     $logMsg\leftarrow message.timestamp$ + $logMsg$
54:     $curationLog\leftarrow curationLog$ + $logMsg$
55:end for
56:return $curationLog$
@@ 4.4 Software and Hardware Artifacts
Our solution has been deployed in a Linux server running Ubuntu 22.04 as the Operating System (OS). The server is equipped with a 6-core 2.6 GHz CPU, 64 GB of memory, and a 4 GB GPU.
@@
The experimentation was executed entirely in Gazebo, an open-source solution for 3D robotics simulations. To provide a realistic scenario, we used a hospital simulation environment made available by Amazon Web Services (AWS) [^2] https://github.com/aws-robotics/aws-robomaker-hospital-world. Additionally, we selected RB-1, an autonomous mobile robot based on the ROS 2 platform produced by Robotnik [69]. This combination of tools and platforms enabled us to assess the impact and effectiveness of our approach.
@@
Table 2: Robag topics rate and average rate per message
@@ 5 Evaluation
The experiments conducted within this proposal focus on navigation tasks included in the Navigation Functionality of the European Robotics League Consumer Service Robots Challenge (ERL Consumer) [74].
@@ 5.1 Black Box Recorder Evaluation
Therefore, our fifth recording method is designed to store a hash for every message within a topic with a rate below 0.5 Hz, one hash for every 5 messages for topics with rates below 1 Hz, a hash for every 10 messages if the rate is below 3 Hz, one hash for every 15 messages if the topic’s rate is below 25 Hz, a hash for every 50 messages for topics with rates below 45 Hz, and one hash for every 100 and 1000 messages when the rate falls below 100 Hz and 400 Hz, respectively. This approach, together with those operating at fixed intervals, has been compared with the generation of a Rosbag file that covers identical topics but lacks anti-tampering properties. Each experiment has been repeated six times, totalling 108 runs, to derive insights into the characteristics of the black box-like device.
@@ 5.1.1 Metrics
Next, regarding system performance, we studied variables such as CPU, memory and disk usage. Specifically, we examined the following variables: system load 1-, 5- and 15-minute average, CPU usage (%), memory usage (GB), sent network traffic (MB/s) and disk writing rate (MB/s). These variables are highly significant for system-overhead analysis given that elevated load conditions can potentially influence the robot’s behavior and event-recording capabilities. In particular, we have obtained a full report about system load, CPU, RAM, network traffic and disk write operations updated every second, facilitating the tracking of overall system activity and an easy comparison with preceding values. To describe the overall distribution and characteristics of these results, we calculated their average as a statistical measure. This information is included in Section 6.
@@ 5.2 Explainability Evaluation
Despite the remarkable performance of LLMs across diverse language tasks, evaluating these models presents challenges, especially in ensuring factual consistency. The extensive capabilities of LLMs and the limitations of current benchmarks in capturing human preferences have given rise to emerging trends in LLM evaluations, such as employing strong LLMs as judges [59]. Traditional similarity-based metrics, such as BLEU and ROUGE, have proven to misalign with human evaluation. While human evaluation remains the primary standard for assessing human preferences, it is exceptionally time-consuming and expensive. To automate the evaluation process, several works have adopted the ’LLM-as-a-judge’ approach, which focuses on using strong LLMs like GPT-4. These models are often trained with RLHF, resulting in strong human alignment [75]. Previous studies have evidenced the consistent agreement of these models with human grading in over 80% of judgments [76, 77, 78]. Given these characteristics, we opted for GPT-4 to assess the answers generated by our explainability component. Through LLM-as-a-judge, we can obtain not only scores but also explanations, making their outputs interpretable. The whole process is represented in Figure 5.
@@
Therefore, to assess the performance of the selected LLM on specific attributes, our approach also includes a criteria-based evaluation covering accuracy, conciseness, coherence, and helpfulness. Selected criteria metrics are described in Table 3. This structured assessment offers a means to assess the model’s effectiveness in generating responses, providing a broader understanding of the model’s capabilities and its alignment with the used scenarios. Employing a scorecard defined for each of these features, our solution assigns scores between 0 and 10 to the responses it generates. The use of this scorecard facilitates a detailed and quantitative analysis of the LLM’s output, ensuring a comprehensive understanding of both its strengths and areas for improvement. A complete example of the answers and evaluations obtained for a selected question in the three selected scenarios is included in Figure 6.
Table 3: Criteria-based evaluation metrics
Criteria metric Description
@@
Figure 6: Examples of correct answers, RAG, and criteria-based evaluations obtained across the three scenarios for the question ’How many goals have been reached by the robot?’
@@ 5.2.1 Categories and Evaluation Questions
To evaluate our explainability proposal in the context of ROS 2 messages interpretation, we have defined a set of questions covering diverse aspects of the navigation process over the how, the what, the which, the when and the where dimensions. These questions provide insights into the different stages and components of the three scenarios. To structure and improve the understanding and analysis of this set of questions, we have organized them into five different categories. These categories range from high-level overviews of navigation to detailed inquiries about trajectory planning, goal completion, and the functioning of the behavior tree. The categories also delve into the causes of unexpected behavior or errors in navigation, improving the explainability of the system whenever an unexpected action occurs. Through these questions, we aim to address key concerns in XAI, including understanding why decisions are made, establishing causality between actions, determining success or failure, identifying timing, and disguising the methods by which achievements are realized [79].

### v3

@@ 3.1.2 Smart Contract
Calling an SC function intended for blockchain writing involves the execution of a transaction that changes the state of the storage data on the blockchain. In our approach, this transaction is signed with the account’s private key via the Elliptic Curve Digital Signature Algorithm, effectively identifying the signer.
@@ 3.2 Explainability Component
Effective communication is essential for achieving explainability in autonomous robots. Our approach is focused on improving this feature through the use of NLP technology, enabling non-expert users to interact with agents in a meaningful way.
@@ 4.1 Black Box Recorder
ROS is the most widely used framework in robotics, offering a collection of software libraries and tools for developing robot applications. Nevertheless, limitations in fulfilling the requirements of real-time embedded systems and the necessity to enhance multi-platform support led to a major upgrade to ROS 2, whose main change lies in the adoption of the Data Distribution Service (DDS) [83].
Nodes constitute the fundamental components in both ROS and ROS 2 applications. These nodes are individual processes that handle specific tasks and communicate with each other through messages. Data transfer between nodes can be achieved using various methods, such as topics, services, and actions. Topics facilitate message passing, services enable synchronous request-reply interactions, and actions are utilized for long-running tasks that need feedback.
This work focuses on navigation capabilities. For this purpose, we utilized Nav2, a suite of tools for ROS 2 that is regarded as the successor to the ROS Navigation Stack. Nav2 equips the robot with the capacity to execute complex navigation tasks.
Nav2’s core functionalities include planning, control, localization, visualization, and additional features. Furthermore, it employs BTs to create customized navigation behavior by orchestrating multiple independent modular servers, offering high flexibility in navigation tasks and the specification of complex robot behaviors [84]. Different independent task servers, intended to compute a path, control effort, or recovery, are communicated with the BT over a ROS 2 interface, such as an action server or service. The use of BTs to generate explanations for robot behavior has been previously successfully evaluated, allowing seeking causal information to answer questions [85].
Within ROS 2, nodes communicate by publishing data on topics, which can be consumed by other nodes. The data itself can range from sensor readings that provide the robot’s perception of its environment to control commands sent to actuators from decision-making nodes. Navigation reproducibility of autonomous robots in ROS 2 involves storing essential data, such as their poses, maps, odometry, velocity, and laser scan readings. To this aim, we have used Rosbag files, designed to capture and store data generated during the execution of a ROS system. These files can reproduce the exact conditions and inputs from a previous run, making them particularly valuable for understanding system behavior. The set of topics included in the Rosbag files generated are detailed in Table 1.
@@
To ensure the immutability of generated Rosbag files, Ethereum has been selected as the blockchain framework. Ethereum, an open-source public blockchain, introduces the cryptocurrency Ether and self-executing programs (SCs) for code execution on the blockchain. Ethereum’s open-source nature ensures decentralization and accessibility, allowing participation without needing approval, a feature crucial for maintaining the integrity of recorded data. Additionally, Ethereum’s use of PoS as a consensus protocol enhances energy efficiency and throughput compared to traditional protocols such as Proof of Work. In addition, Ethereum’s scalability, security and economic efficiency make it a very secure option for ensuring data immutability. The high-level Algorithms 1 and 2 depict the main steps in recording and proof verification services.
@@ 4.2 Log Curation Component
The ROS 2 NavigateToPose action server commands the robot to navigate to a specific goal. While Nav2 usually prioritizes the shortest path, external factors may require adjustments. The /plan topic provides the sequence of poses planned to reach the goal. If unexpected circumstances cause a deviation from this path, the Euclidean distance between consecutive poses in the precomputed plan will increase. This change, together with sensor data from the /scan topic, can indicate potential obstacles causing re-planning. These reasonings have been successfully checked in previous authors’ work [86] to build explanations for autonomous agents based on the use of ROS 2 topics.
@@
6:     if $msgTopic$ == ’/plan’  then
7:         $logMsg\leftarrow$ Log significant path changes and obstacles.
8:     else if $msgTopic$ == ’/navigate_to_pose/_action/status’  then
9:         $logMsg\leftarrow$ Track navigation status and log outcomes.
10:     else if $msgTopic$ == ’/behavior_tree_log’  then
11:         $logMsg\leftarrow$ Log behavior tree node status.
12:     else if $msgTopic$ == ’/amcl_pose’  then
13:         $logMsg\leftarrow$ Record robot position and orientation.
14:     else if $msgTopic$ == ’/cmd_vel’  then
15:         $logMsg\leftarrow$ Log linear and angular velocities.
@@
17:     Append time-stamped $logMsg$ to $curationLog$
18:end for
19:return $curationLog$
@@ 4.4 Software and Hardware Artifacts
Our solution has been deployed in a Linux server running Ubuntu 22.04 as the operating system. The server is equipped with a 6-core 2.6 GHz CPU, 64 GB of memory, and a 4 GB GPU.
@@
The experimentation was executed entirely in Gazebo, an open-source solution for 3D robotics simulations. To provide a realistic scenario, we used a hospital simulation environment made available by Amazon Web Services [^2] https://github.com/aws-robotics/aws-robomaker-hospital-world. Additionally, we selected RB-1, an autonomous mobile robot based on the ROS 2 platform produced by Robotnik [87]. This combination of tools and platforms enabled us to assess the impact and effectiveness of our approach.
@@
Table 2: Rosbag topics rate and average rate per message.
@@ 5 Evaluation
The experiments conducted within this proposal focus on navigation tasks included in the Navigation Functionality of the European Robotics League Consumer Service Robots Challenge [92].
@@ 5.1 Black Box Recorder Evaluation
Therefore, our fifth recording method is designed to store a hash for every message within a topic with a rate below 0.5 Hz, one hash for every 5 messages for topics with rates below 1 Hz, a hash for every 10 messages if the rate is below 3 Hz, one hash for every 15 messages if the topic’s rate is below 25 Hz, a hash for every 50 messages for topics with rates below 45 Hz, and one hash for every 100 and 1000 messages when the rate falls below 100 Hz and 400 Hz, respectively. This approach, together with those operating at fixed intervals, has been compared with the generation of a Rosbag file that covers identical topics but lacks anti-tampering properties. Each experiment has been repeated six times, totaling 108 runs, to derive insights into the characteristics of the black box-like device.
@@ 5.1.1 Metrics
Next, regarding system performance, we studied variables such as CPU, memory and disk usage. Specifically, we examined the following variables: system load 1-, 5- and 15-minute average, CPU usage (%), memory usage (GB), sent network traffic (MB/s) and disk writing rate (MB/s). These variables are highly significant for system-overhead analysis given that elevated load conditions can potentially influence the robot’s behavior and event-recording capabilities. In particular, we have obtained a full report about system load, CPU, RAM, network traffic, and disk write operations updated every second, facilitating the tracking of overall system activity and an easy comparison with preceding values. To describe the overall distribution and characteristics of these results, we calculated their average as a statistical measure. This information is included in Section 6.
@@ 5.2 Explainability Evaluation
Despite the remarkable performance of LLMs in various language tasks, evaluating these models presents challenges, especially in ensuring factual consistency. The extensive capabilities of LLMs and the limitations of current benchmarks in capturing human preferences have led to emerging trends in LLM evaluations, such as the use of strong LLMs as judges [82].
Traditional similarity-based metrics, such as BLEU and ROUGE, have proven to misalign with human evaluation. While human evaluation remains the primary standard for assessing human preferences, it is exceptionally time-consuming and expensive. Furthermore, this approach significantly impacts the scalability of evaluations, making it necessary to use more effective and unbiased methods [81, 93]. To automate the evaluation process, several works have adopted the ‘LLM-as-a-judge’ approach, which focuses on using strong LLMs like GPT-4. These models are often trained with RLHF, resulting in strong human alignment [94]. Previous studies have evidenced the consistent agreement of these models with human grading [95, 96, 97, 98]. Moreover, research shows that LLMs can replicate most of conclusions from human evaluations, particularly in tasks involving prediction accuracy and explanation helpfulness, even across varied settings and domains [81, 99, 100].
However, while LLMs excel in scalability and efficiency, their strengths can be enhanced by integrating the nuanced understanding and subjective judgments that humans provide, particularly in complex or domain-specific contexts [101]. For example, human evaluators can detect subtle inconsistencies or interpret ambiguous explanations in tasks that require deeper contextual understanding, which an automated method might overlook [48, 102].
To address these limitations, we propose a hybrid evaluation method that combines the scalability of LLM-based evaluations with the depth of human assessments. In this approach, a strong LLM such as GPT-4 is used to evaluate a broad set of model outputs, ensuring cost-effective and consistent initial assessments. Human evaluators then validate a representative subset of these outputs to uncover overlooked biases or misalignments. This approach integrates the efficiency and scalability of automated methods with the reliability of human judgment, particularly for nuanced cases.
By adopting this hybrid methodology, we balance efficiency and accuracy, ensuring that the evaluation process captures both objective metrics and subjective preferences effectively. This dual-layered evaluation enables comprehensive assessments while addressing the inherent limitations of relying exclusively on either LLMs or humans.
5.2.1 LLM-as-a-judge Evaluation
The need for a robust LLM aligned with human preferences led us to select GPT-4 to evaluate the answers produced by our explainability component. Through LLM-as-a-judge, we can obtain not only scores but also explanations, making their outputs interpretable. The whole process is represented in Figure 5.
@@ 5.2.1 LLM-as-a-judge Evaluation
Therefore, to assess the performance of the selected LLM on specific attributes, this approach also includes a criteria-based evaluation covering accuracy, conciseness, coherence, and helpfulness. Selected criteria metrics are described in Table 3. This structured assessment offers a means to assess the model’s effectiveness in generating responses, providing a broader understanding of the model’s capabilities and its alignment with the used scenarios. Employing a scorecard defined for each of these features, our solution assigns scores between 0 and 10 to the responses it generates. The use of this scorecard facilitates a detailed and quantitative analysis of the LLM’s output, ensuring a comprehensive understanding of both its strengths and areas for improvement. A complete example of the answers and evaluations obtained for a selected question in the three selected scenarios is included in Figure 6.
Table 3: Criteria-based evaluation metrics in LLM-as-a-judge evaluation.
Metric Description
@@
Figure 6: Examples of correct answers, RAG, and criteria-based evaluations obtained across the three scenarios for the question ’How many goals have been reached by the robot?’ with LLM-as-a-judge evaluation.
5.2.2 Human Evaluation
To validate the explanation results generated by the GPT-4 model, we conducted an online study with 17 participants with diverse demographics. The age distribution ranged from 18 to 55 years, with the majority (47.06%) in the 26-35 age group, followed by 35.29% in the 18-25 age group, 11.76% in the 36-45 age group, and 5.88% in the 46-55 age group. In terms of gender, 62.5% were male and 37.5% were female. Participants’ highest level of education varied, with 37.5% holding a PhD, 37.5% having a Master’s Degree, and 25% possessing a Bachelor’s Degree. Regarding expertise in explainability for autonomous agents, 37.5% were beginners, 31.25% had intermediate expertise, 12.5% were advanced, and 18.75% had no experience.
The evaluation process involved three distinct questionnaires, one corresponding to each scenario. Participants assessed 16 explanations per scenario, resulting in a total of 48 explanations evaluated by each individual. Each explanation was rated using five questions derived from the metrics outlined in Table 4, based on the Explanation Satisfaction Scale proposed by Hoffman et al. in [102]. This scale represents a widely recognized framework for measuring human responses to XAI systems. To ensure the evaluation aligned with human preferences, we used a 5-point Likert scale. Answers in this approach are distributed in 1 (I disagree strongly), 2 (I disagree somewhat), 3 (I’m neutral about it), 4 (I agree somewhat), and 5 (I agree strongly).
In addition to the explanation evaluations and optional demographic information, the questionnaire provided general instructions describing the robot’s task and included a map for each scenario. The questionnaires were distributed via a mailing list, and participants were informed about the anonymity and research purposes of the study. Consent was obtained upon submission of their answers. Participation was voluntary, which made the process cost-effective but also introduced potential drawbacks, such as longer response times and biases that may arise from factors like age, education level, or familiarity with autonomous system explainability approaches.
Statistical analysis is essential in human evaluations to ensure rigor and reliability, particularly when interpreting subjective judgments and addressing variability among participants. Likert-scale data, which is ordinal, requires careful statistical treatment to identify significant trends and differences while avoiding misinterpretation of the results. By applying appropriate statistical methods, we strengthen the credibility of our conclusions and provide a robust evaluation of the LLM’s results across scenarios.
To ensure the reliability and robustness of the evaluation, a series of statistical tests were applied. Reliability analysis of the scale criteria was conducted using Cronbach’s Alpha, a robust metric for assessing internal consistency and ensuring that the scale reliably measures the intended constructs, with a commonly accepted threshold for reliability being $\alpha>0.7$ [103, 104].
To further explore the relationships among variables in the dataset, correlation analysis was performed using Pearson correlation matrices, which provide insights into the strength and direction of linear relationships between variables [105].
For inferential analysis, a normality check was conducted using the Shapiro-Wilk test to determine whether the data met the assumptions required for parametric tests [106]. Since the data violated these assumptions, the non-parametric Friedman test was employed [107]. The Friedman test is particularly suited for analyzing repeated measures on the same participants, as it does not require data to follow a normal distribution [108]. This test assessed whether significant differences existed in participant responses across the three scenarios, offering a reliable framework for understanding the impact of different conditions on the perceived quality and effectiveness of explanations.
Finally, post-hoc analysis using the Nemenyi test was performed to examine participant behavior in specific explanations where significant differences were observed, offering a deeper understanding of the underlying patterns in their responses. The statistical analyses were developed through Python scripts (version 3.10.12), utilizing libraries such as pingouin (version 0.5.5) for Cronbach’s Alpha, scipy.stats (version 1.14.1) for the Shapiro-Wilk and Friedman tests, and pandas (version 2.3.3) for data manipulation. This approach ensured accuracy and reproducibility in the execution of the tests.
Table 4: Questions and metrics used for human evaluation.
Metric Question
Understandability From the explanation, I understand how the task proceeded.
Satisfaction The explanation is satisfactory.
Informativeness The explanation has sufficient detail.
Completeness The explanation seems complete.
Usefulness The explanation is useful to my goals
@@ 5.2.3 Categories and Evaluation Questions
To evaluate our explainability proposal in the context of ROS 2 messages interpretation, we defined a set of questions covering diverse aspects of the navigation process over the how, the what, the which, the when and the where dimensions. These questions provide insights into the different stages and components of the three scenarios. To structure and improve the understanding and analysis of this set of questions, we organized them into five different categories. These categories range from high-level overviews of navigation to detailed inquiries about trajectory planning, goal completion, and the functioning of the BT. The categories also delve into the causes of unexpected behavior or errors in navigation, improving the explainability of the system whenever an unexpected action occurs. Through these questions, we aim to address key concerns in XAI, including understanding why decisions are made, establishing causality between actions, determining success or failure, identifying timing, and disguising the methods by which achievements are realized [109].